A student threw a ball horizontally out a window 8.0m above the ground. It was caught by another student who was 10.0m horizontally away from a spot directly below the first student.What was the velocity of the ball
1
Expert's answer
2019-11-14T08:58:59-0500

The horizontal distance traveled by a ball


"d=v_0t"

The time of a ball free falling


"t=\sqrt{2h/g}"

So, the initial velocity of a ball


"v_0=\frac{d}{t}=\frac{d}{\sqrt{2h/g}}"

"=\frac{10.0\:\rm m}{\sqrt{2\times 8.0\:\rm m/(9.8\:\rm m/s^2)}}=7.83\:\rm m/s"